OpenStudy (anonymous):

The function f(x)= cot(3x-pi/2) has ___.? A. period pi/3 and asymptote at x=pi/6 B. period 2pi/3 and asymptote at x=pi/6 C. period 2pi/3 and asymptote at x=0 D. period pi/3 and asymptote at x=0

jimthompson5910 (jim_thompson5910):

the period of the general cotangent function shown below is T = pi/B y = A*cot(Bx - C) + D

jimthompson5910 (jim_thompson5910):

let me know if that's enough to get you started

OpenStudy (anonymous):

I still don't quiet understand

jimthompson5910 (jim_thompson5910):

do you see how the Bx and 3x match up?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

so B = 3 and you plug that into T = pi/B to get pi/3 as the period

OpenStudy (anonymous):

oh okay. And then how would I find the asymptote ?

jimthompson5910 (jim_thompson5910):

what happens when you plug in x = 0 ?

OpenStudy (anonymous):

you get 0?

jimthompson5910 (jim_thompson5910):

yes, so because x = 0 produces a single number as an output, this means x = 0 is NOT a vertical asymptote

jimthompson5910 (jim_thompson5910):

that leaves x = pi/6 to be an asymptote sure enough if you plug in x = pi/6, you should get NAN or UND or whatever error your calculator will spit out

OpenStudy (anonymous):

so it's not D?

jimthompson5910 (jim_thompson5910):

it's A actually

jimthompson5910 (jim_thompson5910):

yeah D is out because f(0) is defined

OpenStudy (anonymous):

Okay i understand thank you so much:)

jimthompson5910 (jim_thompson5910):

glad to be of help